OpenStudy (anonymous):

How does solving algebraic equations compare to solving trigonometric equations?

OpenStudy (anonymous):

They both are solved using algebra?

OpenStudy (anonymous):

Do you want a contrast in addition to the comparison? Trigonometric functions are periodic, so it's possible to have an infinite set of periodic solutions.

OpenStudy (anonymous):

 you know c^2 = a^2 + b^2 you know sin theta = b/c and cos theta =a/c you (perhaps) know that sin^2 theta + cos^2 theta = 1 From this you can see that they are often 2 sides of the same coin

OpenStudy (anonymous):

Oops got this backwards you know sin theta = a/c and cos theta =b/c